# Possible molecules of triple-heavy pentaquarks within the extended local hidden gauge formalism

### v1

@@ III Results
Channels $\eta\Omega_{bcc}$ $\eta^{{}^{\prime}}\Omega_{bcc}$ $D\Xi_{bc}$ $D\Xi_{bc}^{{}^{\prime}}$ $\bar{B}\Xi_{cc}$ $D_{s}\Omega_{bc}$ $D_{s}\Omega_{bc}^{{}^{\prime}}$ $\bar{B}_{s}\Omega_{cc}$
$0(\frac{1}{2}^{-})$ $8764.63$ (+++++++)(-+++++++)( - + + + + + + + )
$|g_{i}|$ $0.00$ $0.00$ 1.57 $0.00$ $0.00$ $1.52$ $0.00$ $0.00$
$|1-Z|_{i}$ $0.00$ $0.00$ 0.78 $0.00$ $0.00$ $0.22$ $0.00$ $0.00$
$8768.18-12.71i$ (+++++++)(-+++++++)( - + + + + + + + )
$|g_{i}|$ $0.59$ $0.06$ $0.00$ 2.00 $0.48$ $0.00$ $0.94$ $0.25$
$|1-Z|_{i}$ $0.03$ $0.00$ $0.00$ 0.88 $0.02$ $0.00$ $0.08$ $0.003$
$8875.63-2.27i$ (+++++)(-+--++++)( - + - - + + + + )
$|g_{i}|$ $0.21$ $0.06$ $0.00$ $0.10$ 2.15 $0.00$ $0.51$ $1.29$
$|1-Z|_{i}$ $0.004$ $0.001$ $0.00$ $0.002$ 0.88 $0.00$ $0.03$ $0.10$
$8983.70-42.36i$ (++)(------++)( - - - - - - + + )
$|g_{i}|$ $0.82$ $0.35$ $0.00$ $0.28$ $0.22$ $0.00$ 1.83 $1.41$
$|1-Z|_{i}$ $0.05$ $0.03$ $0.00$ $0.01$ $0.01$ $0.00$ 0.72 $0.17$
Channels $\eta\Omega_{bcc}^{*}$ $\eta^{{}^{\prime}}\Omega_{bcc}^{*}$ $D\Xi_{bc}^{*}$ $\bar{B}\Xi_{cc}^{*}$ $D_{s}\Omega_{bc}^{*}$ $\bar{B}_{s}\Omega_{cc}^{*}$
$0(\frac{3}{2}^{-})$ $8793.07-12.20i$ (+++++)(-+++++)( - + + + + + )
$|g_{i}|$ $0.57$ $0.06$ 2.00 $0.43$ $0.95$ $0.27$
$|1-Z|_{i}$ $0.03$ $0.00$ 0.88 $0.01$ $0.08$ $0.003$
$8928.31-3.65i$ (++++)(-+-$+$)( - + - + + + )
$|g_{i}|$ $0.27$ $0.08$ $0.06$ 2.15 $0.62$ $1.23$
$|1-Z|_{i}$ $0.01$ $0.001$ $0.001$ 0.88 $0.06$ $0.09$
$9009.05-43.28i$ (++)(----++)( - - - - + + )
$|g_{i}|$ $0.82$ $0.35$ $0.31$ $0.31$ 1.80 $1.46$
$|1-Z|_{i}$ $0.05$ $0.03$ $0.01$ $0.01$ 0.72 $0.16$
Channels $\omega\Omega_{bcc}$ $\phi\Omega_{bcc}$ $D^{*}\Xi_{bc}$ $D^{*}\Xi_{bc}^{{}^{\prime}}$ $\bar{B}^{*}\Xi_{cc}$ $D_{s}^{*}\Omega_{bc}$ $D_{s}^{*}\Omega_{bc}^{{}^{\prime}}$ $\bar{B}_{s}^{*}\Omega_{cc}$
$0(\frac{1}{2}^{-},\frac{3}{2}^{-})$ $8909.83$ (+++++++)(-+++++++)( - + + + + + + + )
$|g_{i}|$ $0.00$ $0.00$ $1.49$ $0.00$ $0.00$ 1.51 $0.00$ $0.00$
$|1-Z|_{i}$ $0.00$ $0.00$ 0.78 $0.00$ $0.00$ $0.22$ $0.00$ $0.00$
$8915.24-10.17i$ (+++++++)(-+++++++)( - + + + + + + + )
$|g_{i}|$ $0.58$ $0.58$ $0.00$ 1.65 $1.23$ $0.00$ $0.92$ $0.61$
$|1-Z|_{i}$ $0.04$ $0.04$ $0.00$ 0.65 $0.25$ $0.00$ $0.07$ $0.02$
$8927.25-0.73i$ (+++++++)(-+++++++)( - + + + + + + + )
$|g_{i}|$ $0.15$ $0.13$ $0.00$ $0.88$ 1.76 $0.00$ $0.68$ $1.32$
$|1-Z|_{i}$ $0.003$ $0.003$ $0.00$ $0.23$ 0.69 $0.00$ $0.04$ $0.10$
$9115.00-33.41i$ ($+$)(-----$+$)( - - - - - + + + )
$|g_{i}|$ $0.48$ $0.78$ $0.00$ $0.20$ $0.21$ $0.00$ 1.76 $1.18$
$|1-Z|_{i}$ $0.02$ $0.10$ $0.00$ $0.01$ $0.003$ $0.00$ 0.65 $0.22$
--Continue.--
--Continue.--
Channels $\omega\Omega_{bcc}^{*}$ $\phi\Omega_{bcc}^{*}$ $D^{*}\Xi_{bc}^{*}$ $\bar{B}^{*}\Xi_{cc}^{*}$ $D_{s}^{*}\Omega_{bc}^{*}$ $\bar{B}_{s}^{*}\Omega_{cc}^{*}$
$0(\frac{1}{2}^{-},\frac{3}{2}^{-},\frac{5}{2}^{-})$ $8942.49-10.64i$ (+++++)(-+++++)( - + + + + + )
$|g_{i}|$ $0.57$ $0.55$ 1.79 $0.60$ $1.08$ $0.08$
$|1-Z|_{i}$ $0.04$ $0.04$ 0.79 $0.04$ $0.11$ $0.00$
$8977.58-0.27i$ (+++++)(-+++++)( - + + + + + )
$|g_{i}|$ $0.09$ $0.11$ $0.24$ 2.00 $0.40$ $1.32$
$|1-Z|_{i}$ $0.001$ $0.002$ $0.05$ 0.84 $0.02$ $0.10$
$9141.48-35.55i$ (++)(----++)( - - - - + + )
$|g_{i}|$ $0.49$ $0.78$ $0.23$ $0.13$ 1.75 $1.21$
$|1-Z|_{i}$ $0.02$ $0.09$ $0.01$ $0.002$ 0.68 $0.18$
@@
For the case $VB(\frac{1}{2}^{+})$ degenerate in $J^{P}=\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, we also obtain four states, where the most relevant channels are the $D^{*}\Xi_{bc}$ for the one $8909.83$ MeV, the $D^{*}\Xi_{bc}^{{}^{\prime}}$ for the state $(8915.24-10.17i)$ MeV, the $\bar{B}^{*}\Xi_{cc}$ for the pole $(8927.25-0.73i)$ MeV, and the $D_{s}^{*}\Omega_{bc}^{{}^{\prime}}$ for the last one $(9115.00-33.41i)$ MeV, accordingly. Their binding energies are about $21$ MeV, $41$ MeV, $19$ MeV, and $44$ MeV, respectively. Note that, for the state $D^{*}\Xi_{bc}$ with the mass $8909.83$ MeV, its coupling to the channel $D^{*}\Xi_{bc}$ is $g=1.49$, which is not the biggest one and a bit smaller than the one to the channel $D_{s}^{*}\Omega_{bc}$ of $g=1.51$. Its mass $8909.83$ MeV is about $213$ MeV below the threshold of the $D_{s}^{*}\Omega_{bc}$ channel, where the binding energy for the $D_{s}^{*}\Omega_{bc}$ channel is too large with respect to the others’. But, for the results of the compositenesses, the biggest one is 0.78 for the channel $D^{*}\Xi_{bc}$. In view of the results for the couplings and the compositenesses, the pole $8909.83$ MeV should be mainly contributed from the channel $D^{*}\Xi_{bc}$. Therefore, we make a further checking and investigate the interaction of the single channel, where we find that one pole of $D^{*}\Xi_{bc}$ is bound by $65$ MeV, and the other one of $D_{s}^{*}\Omega_{bc}$ is bound by $41$ MeV. This test result indicates that the pole $8909.83$ MeV found in the coupled channel interaction is dominant from the channel $D^{*}\Xi_{bc}$. Furthermore, if we take the free parameter $\mu=1000$ MeV, the mass of this state becomes $8867.35$ MeV, and its couplings to the two channels are reversed, which are $g=1.94$ for the channel $D^{*}\Xi_{bc}$, and $g=1.79$ for the channel $D_{s}^{*}\Omega_{bc}$. Thus, in this case, one can easily see that the pole $8867.35$ MeV mostly couple to the channel $D^{*}\Xi_{bc}$ with the binding energy about $63$ MeV, which also confirm our conclusion above. Indeed, the channel $D^{*}\Xi_{bc}$ couples strongly to the $D_{s}^{*}\Omega_{bc}$, which leads to the large coupling $g=1.79$ for the pole $8867.35$ MeV and can be seen from the non-diagonal coefficient $\sqrt{2}$ between them in Table 3. Note that, from the results of the coupled channel interaction as shown in Table 6, the pole of the channel $D_{s}^{*}\Omega_{bc}$ is missing, which can be expected from its attractive potential as given in Table 3. In fact, the channel $D_{s}^{*}\Omega_{bc}$ is loose bound with the binding energy about $41$ MeV in the single channel interaction obtained above, which is smaller than the one $65$ MeV for the channel $D^{*}\Xi_{bc}$. Therefore, taking the parameter $\mu=800$ MeV for all the results at present, its pole has moved to the threshold and becomes no stable pole to be found, which is similar situation for the other systems with the states $\Omega_{cc}^{(*)}$, $\Omega_{bb}^{(*)}$, some of the systems with the $\Omega_{bc}^{(*)}$, and some of the cases with the $\Omega_{bc}^{\prime}$.
@@
Channels $\eta\Omega_{bbc}$ $\eta^{{}^{\prime}}\Omega_{bbc}$ $D\Xi_{bb}$ $\bar{B}\Xi_{bc}$ $\bar{B}\Xi_{bc}^{{}^{\prime}}$ $D_{s}\Omega_{bb}$ $\bar{B}_{s}\Omega_{bc}$ $\bar{B}_{s}\Omega_{bc}^{{}^{\prime}}$
$0(\frac{1}{2}^{-})$ $12134.09-0.01i$ (+++++++)(-+++++++)( - + + + + + + + )
$|g_{i}|$ $0.01$ $0.14$ 1.75 $0.01$ $0.01$ $1.28$ $0.004$ $0.003$
$|1-Z|_{i}$ $0.00$ $0.01$ 0.62 $0.00$ $0.00$ $0.36$ $0.00$ $0.00$
$12189.43-7.63i$ (++++++)(--++++++)( - - + + + + + + )
$|g_{i}|$ $0.44$ $0.13$ $0.43$ 1.59 $0.37$ $0.49$ $1.05$ $0.05$
$|1-Z|_{i}$ $0.02$ $0.005$ $0.08$ 1.10 $0.04$ $0.13$ $0.11$ $0.00$
$12217.65-1.30i$ ($+$)(----+-++)( - - - - + - + + )
$|g_{i}|$ $0.11$ $0.03$ $0.10$ $0.08$ 1.32 $0.14$ $0.15$ $0.93$
$|1-Z|_{i}$ $0.001$ $0.00$ $0.01$ $0.002$ 0.91 $0.01$ $0.003$ $0.09$
Channels $\eta\Omega_{bbc}^{*}$ $\eta^{{}^{\prime}}\Omega_{bbc}^{*}$ $D\Xi_{bb}^{*}$ $\bar{B}\Xi_{bc}^{*}$ $D_{s}\Omega_{bb}^{*}$ $\bar{B}_{s}\Omega_{bc}^{*}$
$0(\frac{3}{2}^{-})$ $12163.34$ (+++++)(-+++++)( - + + + + + )
$|g_{i}|$ $0.01$ $0.14$ 1.74 $0.01$ $1.29$ $0.01$
$|1-Z|_{i}$ $0.00$ $0.01$ 0.61 $0.00$ $0.37$ $0.00$
$12243.89-5.08i$ (++)(---+-+)( - - - + - + )
$|g_{i}|$ $0.23$ $0.07$ $0.21$ 1.27 $0.29$ $0.98$
$|1-Z|_{i}$ $0.004$ $0.001$ $0.03$ 0.86 $0.04$ $0.10$
Channels $\omega\Omega_{bbc}$ $\phi\Omega_{bbc}$ $D^{*}\Xi_{bb}$ $\bar{B}^{*}\Xi_{bc}$ $\bar{B}^{*}\Xi_{bc}^{{}^{\prime}}$ $D_{s}^{*}\Omega_{bb}$ $\bar{B}_{s}^{*}\Omega_{bc}$ $\bar{B}_{s}^{*}\Omega_{bc}^{{}^{\prime}}$
$0(\frac{1}{2}^{-},\frac{3}{2}^{-})$ $12233.59-1.54i$ (++++++)(--++++++)( - - + + + + + + )
$|g_{i}|$ $0.16$ $0.18$ $0.15$ 1.42 $0.13$ $0.18$ $0.87$ $0.01$
$|1-Z|_{i}$ $0.002$ $0.01$ $0.004$ 0.91 $0.004$ $0.01$ $0.08$ $0.00$
$12262.25-0.63i$ (+++++)(--+-++++)( - - + - + + + + )
$|g_{i}|$ $0.08$ $0.10$ $0.10$ $0.06$ 1.34 $0.10$ $0.12$ $0.89$
$|1-Z|_{i}$ $0.001$ $0.003$ $0.002$ $0.002$ 0.92 $0.002$ $0.002$ $0.08$
$12278.80-0.66i$ (++++)(--+--$+$)( - - + - - + + + )
$|g_{i}|$ $0.09$ $0.09$ 1.72 $0.01$ $0.01$ $1.27$ $0.01$ $0.01$
$|1-Z|_{i}$ $0.001$ $0.002$ 0.63 $0.00$ $0.00$ $0.36$ $0.00$ $0.00$
Channels $\omega\Omega_{bbc}^{*}$ $\phi\Omega_{bbc}^{*}$ $D^{*}\Xi_{bb}^{*}$ $\bar{B}^{*}\Xi_{bc}^{*}$ $D_{s}^{*}\Omega_{bb}^{*}$ $\bar{B}_{s}^{*}\Omega_{bc}^{*}$
$0(\frac{1}{2}^{-},\frac{3}{2}^{-},\frac{5}{2}^{-})$ $12283.84-2.78i$ (++++)(--++++)( - - + + + + )
$|g_{i}|$ $0.19$ $0.23$ $0.20$ 1.45 $0.24$ $0.83$
$|1-Z|_{i}$ $0.003$ $0.01$ $0.01$ 0.92 $0.01$ $0.07$
$12308.10-0.68i$ ($+$)(--+-++)( - - + - + + )
$|g_{i}|$ $0.09$ $0.10$ 1.72 $0.01$ $1.27$ $0.02$
$|1-Z|_{i}$ $0.001$ $0.002$ 0.62 $0.00$ $0.37$ $0.00$
@@
In Table 7, we show the poles and their couplings, compositenesses for each channel in the $\Omega_{bbc}$-like sector. The results are still obtained by taking $\mu=800$ MeV for the loop functions. We find a total of ten bound states here. For the case $PB(\frac{1}{2}^{+})$ with $J^{P}=\frac{1}{2}^{-}$, the three states are the possible candidates of the molecules $D\Xi_{bb}$, $\bar{B}\Xi_{bc}$, and $\bar{B}\Xi_{bc}^{{}^{\prime}}$, respectively, in view of the results of the couplings and compositenesses. For the case $PB(\frac{3}{2}^{+})$ with $J^{P}=\frac{3}{2}^{-}$, the two states are the possible molecules $D\Xi_{bb}^{*}$ and $\bar{B}\Xi_{bc}^{*}$, respectively. For the case $VB(\frac{1}{2}^{+})$ degenerate in $J^{P}=\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, three possible molecular states are found in the channels $\bar{B}^{*}\Xi_{bc}$, $\bar{B}^{*}\Xi_{bc}^{{}^{\prime}}$, and $D^{*}\Xi_{bb}$, respectively. For the case $VB(\frac{3}{2}^{+})$ degenerate in $J^{P}=\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, $\frac{5}{2}^{-}$, the two bound systems are the $\bar{B}^{*}\Xi_{bc}^{*}$ and $D^{*}\Xi_{bb}^{*}$ with respective poles. Most of the binding energies are about $10$ MeV more or less, and the ones for the other four are about $70$ MeV, which are larger than those of the $\Omega_{bcc}$ sector.

### v3

@@ III Results
$0(\frac{1}{2}^{-})$ Channels $\eta\Omega_{bcc}$ $\eta^{{}^{\prime}}\Omega_{bcc}$ $D\Xi_{bc}$ $D\Xi_{bc}^{{}^{\prime}}$ $\bar{B}\Xi_{cc}$ $D_{s}\Omega_{bc}$ $D_{s}\Omega_{bc}^{{}^{\prime}}$ $\bar{B}_{s}\Omega_{cc}$
$8764.63$ $|g_{i}|$ $0.00$ $0.00$ 1.57 $0.00$ $0.00$ $1.52$ $0.00$ $0.00$
(+++++++)(-+++++++)( - + + + + + + + ) $|1-Z|_{i}$ $0.00$ $0.00$ 0.78 $0.00$ $0.00$ $0.22$ $0.00$ $0.00$
$8768.18-12.71i$ $|g_{i}|$ $0.59$ $0.06$ $0.00$ 2.00 $0.48$ $0.00$ $0.94$ $0.25$
(+++++++)(-+++++++)( - + + + + + + + ) $|1-Z|_{i}$ $0.03$ $0.00$ $0.00$ 0.88 $0.02$ $0.00$ $0.08$ $0.003$
$8875.63-2.27i$ $|g_{i}|$ $0.21$ $0.06$ $0.00$ $0.10$ 2.15 $0.00$ $0.51$ $1.29$
(+++++)(-+--++++)( - + - - + + + + ) $|1-Z|_{i}$ $0.004$ $0.001$ $0.00$ $0.002$ 0.88 $0.00$ $0.03$ $0.10$
$8983.70-42.36i$ $|g_{i}|$ $0.82$ $0.35$ $0.00$ $0.28$ $0.22$ $0.00$ 1.83 $1.41$
(++)(------++)( - - - - - - + + ) $|1-Z|_{i}$ $0.05$ $0.03$ $0.00$ $0.01$ $0.01$ $0.00$ 0.72 $0.17$
$0(\frac{3}{2}^{-})$ Channels $\eta\Omega_{bcc}^{*}$ $\eta^{{}^{\prime}}\Omega_{bcc}^{*}$ $D\Xi_{bc}^{*}$ $\bar{B}\Xi_{cc}^{*}$ $D_{s}\Omega_{bc}^{*}$ $\bar{B}_{s}\Omega_{cc}^{*}$
$8793.07-12.20i$ $|g_{i}|$ $0.57$ $0.06$ 2.00 $0.43$ $0.95$ $0.27$
(+++++)(-+++++)( - + + + + + ) $|1-Z|_{i}$ $0.03$ $0.00$ 0.88 $0.01$ $0.08$ $0.003$
$8928.31-3.65i$ $|g_{i}|$ $0.27$ $0.08$ $0.06$ 2.15 $0.62$ $1.23$
(++++)(-+-$+$)( - + - + + + ) $|1-Z|_{i}$ $0.01$ $0.001$ $0.001$ 0.88 $0.06$ $0.09$
$9009.05-43.28i$ $|g_{i}|$ $0.82$ $0.35$ $0.31$ $0.31$ 1.80 $1.46$
(++)(----++)( - - - - + + ) $|1-Z|_{i}$ $0.05$ $0.03$ $0.01$ $0.01$ 0.72 $0.16$
$0(\frac{1}{2}^{-},\frac{3}{2}^{-})$ Channels $\omega\Omega_{bcc}$ $\phi\Omega_{bcc}$ $D^{*}\Xi_{bc}$ $D^{*}\Xi_{bc}^{{}^{\prime}}$ $\bar{B}^{*}\Xi_{cc}$ $D_{s}^{*}\Omega_{bc}$ $D_{s}^{*}\Omega_{bc}^{{}^{\prime}}$ $\bar{B}_{s}^{*}\Omega_{cc}$
$8909.83$ $|g_{i}|$ $0.00$ $0.00$ $1.49$ $0.00$ $0.00$ 1.51 $0.00$ $0.00$
(+++++++)(-+++++++)( - + + + + + + + ) $|1-Z|_{i}$ $0.00$ $0.00$ 0.78 $0.00$ $0.00$ $0.22$ $0.00$ $0.00$
$8915.24-10.17i$ $|g_{i}|$ $0.58$ $0.58$ $0.00$ 1.65 $1.23$ $0.00$ $0.92$ $0.61$
(+++++++)(-+++++++)( - + + + + + + + ) $|1-Z|_{i}$ $0.04$ $0.04$ $0.00$ 0.65 $0.25$ $0.00$ $0.07$ $0.02$
$8927.25-0.73i$ $|g_{i}|$ $0.15$ $0.13$ $0.00$ $0.88$ 1.76 $0.00$ $0.68$ $1.32$
(+++++++)(-+++++++)( - + + + + + + + ) $|1-Z|_{i}$ $0.003$ $0.003$ $0.00$ $0.23$ 0.69 $0.00$ $0.04$ $0.10$
$9115.00-33.41i$ $|g_{i}|$ $0.48$ $0.78$ $0.00$ $0.20$ $0.21$ $0.00$ 1.76 $1.18$
($+$)(-----$+$)( - - - - - + + + ) $|1-Z|_{i}$ $0.02$ $0.10$ $0.00$ $0.01$ $0.003$ $0.00$ 0.65 $0.22$
$0(\frac{1}{2}^{-},\frac{3}{2}^{-},\frac{5}{2}^{-})$ Channels $\omega\Omega_{bcc}^{*}$ $\phi\Omega_{bcc}^{*}$ $D^{*}\Xi_{bc}^{*}$ $\bar{B}^{*}\Xi_{cc}^{*}$ $D_{s}^{*}\Omega_{bc}^{*}$ $\bar{B}_{s}^{*}\Omega_{cc}^{*}$
$8942.49-10.64i$ $|g_{i}|$ $0.57$ $0.55$ 1.79 $0.60$ $1.08$ $0.08$
(+++++)(-+++++)( - + + + + + ) $|1-Z|_{i}$ $0.04$ $0.04$ 0.79 $0.04$ $0.11$ $0.00$
$8977.58-0.27i$ $|g_{i}|$ $0.09$ $0.11$ $0.24$ 2.00 $0.40$ $1.32$
(+++++)(-+++++)( - + + + + + ) $|1-Z|_{i}$ $0.001$ $0.002$ $0.05$ 0.84 $0.02$ $0.10$
$9141.48-35.55i$ $|g_{i}|$ $0.49$ $0.78$ $0.23$ $0.13$ 1.75 $1.21$
(++)(----++)( - - - - + + ) $|1-Z|_{i}$ $0.02$ $0.09$ $0.01$ $0.002$ 0.68 $0.18$
@@
For the case $VB(\frac{1}{2}^{+})$ degenerate in $J^{P}=\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, we also obtain four states, where the most relevant channels are the $D^{*}\Xi_{bc}$ for the one $8909.83$ MeV, the $D^{*}\Xi_{bc}^{{}^{\prime}}$ for the state $(8915.24-10.17i)$ MeV, the $\bar{B}^{*}\Xi_{cc}$ for the pole $(8927.25-0.73i)$ MeV, and the $D_{s}^{*}\Omega_{bc}^{{}^{\prime}}$ for the last one $(9115.00-33.41i)$ MeV, accordingly. Their binding energies are about $21$ MeV, $41$ MeV, $19$ MeV, and $44$ MeV, respectively. Note that, for the state $D^{*}\Xi_{bc}$ with the mass $8909.83$ MeV, its coupling to the channel $D^{*}\Xi_{bc}$ is $g=1.49$, which is not the biggest one and a bit smaller than the one to the channel $D_{s}^{*}\Omega_{bc}$ of $g=1.51$. Its mass $8909.83$ MeV is about $213$ MeV below the threshold of the $D_{s}^{*}\Omega_{bc}$ channel, where the binding energy for the $D_{s}^{*}\Omega_{bc}$ channel is too large with respect to the other ones. But, for the results of the compositenesses, the biggest one is 0.78 for the channel $D^{*}\Xi_{bc}$. In view of the results for the couplings and the compositenesses, the pole $8909.83$ MeV should be mainly contributed from the channel $D^{*}\Xi_{bc}$. Therefore, we make a further checking and investigate the interaction of the single channel, where we find that one pole of $D^{*}\Xi_{bc}$ is bound by $65$ MeV, and the other one of $D_{s}^{*}\Omega_{bc}$ is bound by $41$ MeV. This test result indicates that the pole $8909.83$ MeV found in the coupled channel interaction is dominant by the channel $D^{*}\Xi_{bc}$. Furthermore, if we take the free parameter $\mu=1000$ MeV, the mass of this state becomes $8867.35$ MeV, and its couplings to the two channels are reversed, which are $g=1.94$ for the channel $D^{*}\Xi_{bc}$, and $g=1.79$ for the channel $D_{s}^{*}\Omega_{bc}$. Thus, in this case, one can easily see that the pole $8867.35$ MeV mostly couples to the channel $D^{*}\Xi_{bc}$ with the binding energy about $63$ MeV, which also confirm our conclusion above. Indeed, the channel $D^{*}\Xi_{bc}$ couples strongly to the $D_{s}^{*}\Omega_{bc}$, which leads to the large coupling $g=1.79$ for the pole $8867.35$ MeV and can be seen from the non-diagonal coefficient $\sqrt{2}$ between them in Table 3. Note that, from the results of the coupled channel interaction as shown in Table 6, the pole of the channel $D_{s}^{*}\Omega_{bc}$ is missing, which can be expected from its attractive potential as given in Table 3. In fact, the channel $D_{s}^{*}\Omega_{bc}$ is loosely bound with the binding energy about $41$ MeV in the single channel interaction obtained above, which is smaller than the one $65$ MeV for the channel $D^{*}\Xi_{bc}$. Therefore, taking the parameter $\mu=800$ MeV for all the results at present, its pole has moved to the threshold and no stable pole to be found, which is a similar situation for the other systems with the states $\Omega_{cc}^{(*)}$, $\Omega_{bb}^{(*)}$, some of the systems with the $\Omega_{bc}^{(*)}$, and some of the cases with the $\Omega_{bc}^{\prime}$.
@@
$0(\frac{1}{2}^{-})$ Channels $\eta\Omega_{bbc}$ $\eta^{{}^{\prime}}\Omega_{bbc}$ $D\Xi_{bb}$ $\bar{B}\Xi_{bc}$ $\bar{B}\Xi_{bc}^{{}^{\prime}}$ $D_{s}\Omega_{bb}$ $\bar{B}_{s}\Omega_{bc}$ $\bar{B}_{s}\Omega_{bc}^{{}^{\prime}}$
$12134.09-0.01i$ $|g_{i}|$ $0.01$ $0.14$ 1.75 $0.01$ $0.01$ $1.28$ $0.004$ $0.003$
(+++++++)(-+++++++)( - + + + + + + + ) $|1-Z|_{i}$ $0.00$ $0.01$ 0.62 $0.00$ $0.00$ $0.36$ $0.00$ $0.00$
$12189.43-7.63i$ $|g_{i}|$ $0.44$ $0.13$ $0.43$ 1.59 $0.37$ $0.49$ $1.05$ $0.05$
(++++++)(--++++++)( - - + + + + + + ) $|1-Z|_{i}$ $0.02$ $0.005$ $0.08$ 1.10 $0.04$ $0.13$ $0.11$ $0.00$
$12217.65-1.30i$ $|g_{i}|$ $0.11$ $0.03$ $0.10$ $0.08$ 1.32 $0.14$ $0.15$ $0.93$
($+$)(----+-++)( - - - - + - + + ) $|1-Z|_{i}$ $0.001$ $0.00$ $0.01$ $0.002$ 0.91 $0.01$ $0.003$ $0.09$
$0(\frac{3}{2}^{-})$ Channels $\eta\Omega_{bbc}^{*}$ $\eta^{{}^{\prime}}\Omega_{bbc}^{*}$ $D\Xi_{bb}^{*}$ $\bar{B}\Xi_{bc}^{*}$ $D_{s}\Omega_{bb}^{*}$ $\bar{B}_{s}\Omega_{bc}^{*}$
$12163.34$ $|g_{i}|$ $0.01$ $0.14$ 1.74 $0.01$ $1.29$ $0.01$
(+++++)(-+++++)( - + + + + + ) $|1-Z|_{i}$ $0.00$ $0.01$ 0.61 $0.00$ $0.37$ $0.00$
$12243.89-5.08i$ $|g_{i}|$ $0.23$ $0.07$ $0.21$ 1.27 $0.29$ $0.98$
(++)(---+-+)( - - - + - + ) $|1-Z|_{i}$ $0.004$ $0.001$ $0.03$ 0.86 $0.04$ $0.10$
$0(\frac{1}{2}^{-},\frac{3}{2}^{-})$ Channels $\omega\Omega_{bbc}$ $\phi\Omega_{bbc}$ $D^{*}\Xi_{bb}$ $\bar{B}^{*}\Xi_{bc}$ $\bar{B}^{*}\Xi_{bc}^{{}^{\prime}}$ $D_{s}^{*}\Omega_{bb}$ $\bar{B}_{s}^{*}\Omega_{bc}$ $\bar{B}_{s}^{*}\Omega_{bc}^{{}^{\prime}}$
$12233.59-1.54i$ $|g_{i}|$ $0.16$ $0.18$ $0.15$ 1.42 $0.13$ $0.18$ $0.87$ $0.01$
(++++++)(--++++++)( - - + + + + + + ) $|1-Z|_{i}$ $0.002$ $0.01$ $0.004$ 0.91 $0.004$ $0.01$ $0.08$ $0.00$
$12262.25-0.63i$ $|g_{i}|$ $0.08$ $0.10$ $0.10$ $0.06$ 1.34 $0.10$ $0.12$ $0.89$
(+++++)(--+-++++)( - - + - + + + + ) $|1-Z|_{i}$ $0.001$ $0.003$ $0.002$ $0.002$ 0.92 $0.002$ $0.002$ $0.08$
$12278.80-0.66i$ $|g_{i}|$ $0.09$ $0.09$ 1.72 $0.01$ $0.01$ $1.27$ $0.01$ $0.01$
(++++)(--+--$+$)( - - + - - + + + ) $|1-Z|_{i}$ $0.001$ $0.002$ 0.63 $0.00$ $0.00$ $0.36$ $0.00$ $0.00$
$0(\frac{1}{2}^{-},\frac{3}{2}^{-},\frac{5}{2}^{-})$ Channels $\omega\Omega_{bbc}^{*}$ $\phi\Omega_{bbc}^{*}$ $D^{*}\Xi_{bb}^{*}$ $\bar{B}^{*}\Xi_{bc}^{*}$ $D_{s}^{*}\Omega_{bb}^{*}$ $\bar{B}_{s}^{*}\Omega_{bc}^{*}$
$12283.84-2.78i$ $|g_{i}|$ $0.19$ $0.23$ $0.20$ 1.45 $0.24$ $0.83$
(++++)(--++++)( - - + + + + ) $|1-Z|_{i}$ $0.003$ $0.01$ $0.01$ 0.92 $0.01$ $0.07$
$12308.10-0.68i$ $|g_{i}|$ $0.09$ $0.10$ 1.72 $0.01$ $1.27$ $0.02$
($+$)(--+-++)( - - + - + + ) $|1-Z|_{i}$ $0.001$ $0.002$ 0.62 $0.00$ $0.37$ $0.00$
@@
In Table 7, we show the poles and their couplings, and the compositenesses for each channel in the $\Omega_{bbc}$-like sector. The results are still obtained by taking $\mu=800$ MeV for the loop functions. We find a total of ten bound states here. For the case $PB(\frac{1}{2}^{+})$ with $J^{P}=\frac{1}{2}^{-}$, the three states are the possible candidates of the molecules $D\Xi_{bb}$, $\bar{B}\Xi_{bc}$, and $\bar{B}\Xi_{bc}^{{}^{\prime}}$, respectively, in view of the results of the couplings and compositenesses. For the case $PB(\frac{3}{2}^{+})$ with $J^{P}=\frac{3}{2}^{-}$, the two states are the possible molecules $D\Xi_{bb}^{*}$ and $\bar{B}\Xi_{bc}^{*}$, respectively. For the case $VB(\frac{1}{2}^{+})$ degenerate in $J^{P}=\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, three possible molecular states are found in the channels $\bar{B}^{*}\Xi_{bc}$, $\bar{B}^{*}\Xi_{bc}^{{}^{\prime}}$, and $D^{*}\Xi_{bb}$, respectively. For the case $VB(\frac{3}{2}^{+})$ degenerate in $J^{P}=\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, $\frac{5}{2}^{-}$, the two bound systems are the $\bar{B}^{*}\Xi_{bc}^{*}$ and $D^{*}\Xi_{bb}^{*}$ with respective poles. Most of the binding energies are about $10$ MeV more or less, and the ones for the other four are about $70$ MeV, which are larger than those of the $\Omega_{bcc}$ sector.